OpenStudy (anonymous):

Write the equation of the line perpendicular to 4x – 5y = -10 through the point (2, 3). Write the equation in slope intercept form.

OpenStudy (superdavesuper):

re-arrange ur eqn to 5y=4x+10, so wats the slope?

OpenStudy (anonymous):

ignore the last 2 lines

OpenStudy (anonymous):

\[y = \frac{ 4 }{ 5 }x +2\] So the line we are finding must have slope of -5/4 \[y = \frac{ -5 }{ 4 }x +b\] That line goes through (2,3) then \[3 = \frac{ -5 }{ 4 }2 +b = \frac{-10}{4}+b\] So b = 11/2 The line is: \[y = \frac{-5}{4}x+11/2\]

OpenStudy (anonymous):

oh right and simplify 22/4 to 11/2